# Rakuten Data Release: A Large-Scale and Long-Term Reviews Corpus for Hotel Domain

### v3

@@ 1 Introduction
Our data repository, “Rakuten Data Release” [^1] https://rit.rakuten.com/data_release/ (RDR), has enabled academic researchers to undertake data-driven research leveraging consumer data collected across a range of Rakuten services, specifically Rakuten Ichiba [^2] https://www.rakuten.co.jp, Rakuten Travel [^3] https://travel.rakuten.co.jp, Rakuten GORA [^4] https://gora.golf.rakuten.co.jp, and Rakuten Recipe [^5] https://recipe.rakuten.co.jp. In particular, use of Rakuten Service’s customer reviews has brought remarkable success on recommender systems kanezashi:25 and various Natural Language Processing tasks such as word replacement Hiraoka:22, tokenization hiraoka:21, review response generation shirai:24, text normalization higashiyama:25, and sentiment analysis Nakayama:15; toyama:17; Nakayama:22.
@@
Intending to advance and facilitate research using customer review data in the field of Japanese NLP research and the other related fields, in this paper, we provide an up-to-date Rakuten Travel reviews corpus as shown in Table 1, ranging from the pre-COVID era to the post-COVID era. Our main contributions are summarized as follows: (1) To the best of our knowledge, in the hotel domain, no attempt has been made to release a publicly available dataset in which the collection duration is more than 15 years; (2) We distribute our corpus for easy access by any research institute, particularly, university laboratories for academic research purposes, which can be obtained via the IDR repository [^7] https://www.nii.ac.jp/dsc/idr/rakuten/; (3) We give an insight into factors of data drift in hotel review texts with statistical approaches.
@@ 2 Related Work
One of the major review datasets was released by Amazon, Inc. back in 2020 keung:20. As an initiative separate from the Amazon, hou:24 released up-to-date Amazon review datasets named “AMAZON REVIEWS 2023”, ranging from May 1996 to September 2023  [^8] https://amazon-reviews-2023.github.io. There are user-generated review datasets in the hotel domain, mainly crawled from leading online travel platforms such as TripAdvisor and Booking.com. For instance, Fang:16 crawled 41k reviews posted in 2014 for attractions in New Orleans to analyze the perceived helpfulness of reviews. Tsamis:21 crawled 65k English hotel reviews from TripAdvisor and trained a DNN on helpfulness prediction. The Kaggle dataset  [^9] https://www.kaggle.com/datasets/jiashenliu/515k-hotel-reviews-data-in-europe/discussion?sort=undefined contains 515k reviews crawled from Booking.com published between August 2015 and August 2017. Finally, Igebaria:24 created a corpus of English reviews posted in 2023 for predicting individual review helpfulness. Although these datasets are large-scale, they do not provide sufficiently long period, covering three significant eras in the Travel industry: pre-COVID, mid-COVID, and post-COVID. None of these studies focuses on releasing review texts for more than 15 years in the hotel domain.
@@ 4 Factors Causing Data Drift
“エモい” is originated from “emotional”. This is a very convenient and ambiguous term that encompasses a diverse range of emotions and sensations such as "feeling deeply moved," “resonating deeply,“nostalgic,“poignant, and “having a unique atmosphere. Additionally, although there was no significant difference, other new slang words appeared in 2024 whereas in 2019 these did not such as “チルい (chirui) derived from “Chilling out” and “沼る (obsessed)”. Such new slang words make NLP tasks, such as sentiment analysis, more difficult.

### v4

@@ 1 Introduction
Our data repository, “Rakuten Data Release” [^1] https://rit.rakuten.com/data_release/ (RDR), has enabled academic researchers to undertake data-driven research leveraging consumer data collected across a range of Rakuten services, specifically Rakuten Ichiba [^2] https://www.rakuten.co.jp, Rakuten Travel [^3] https://travel.rakuten.co.jp, Rakuten GORA [^4] https://gora.golf.rakuten.co.jp, and Rakuten Recipe [^5] https://recipe.rakuten.co.jp. In particular, use of Rakuten Service’s customer reviews has brought remarkable success on recommender systems Suzumura et al. (2025) and various Natural Language Processing tasks such as word replacement Hiraoka et al. (2022), tokenization Hiraoka et al. (2021), review response generation Shirai et al. (2024), text normalization Higashiyama and Utiyama (2025), and sentiment analysis Nakayama and Fujii (2015); Toyama et al. (2017); Nakayama et al. (2022).
@@
Intending to advance and facilitate research using customer review data in the field of Japanese NLP research and the other related fields from the above perspectives, in this paper, we provide an up-to-date Rakuten Travel reviews corpus as shown in Table 1, ranging from the pre-COVID era to the post-COVID era. Our main contributions are summarized as follows: (1) To the best of our knowledge, in the hotel domain, no attempt has been made to release a publicly available dataset in which the collection duration is more than 15 years; (2) We distribute our corpus for easy access by any research institute, particularly, university laboratories for academic research purposes, which can be obtained via the IDR repository [^7] https://www.nii.ac.jp/dsc/idr/rakuten/; (3) We give an insight into factors of data drift in hotel review texts with statistical approaches.
@@ 2 Related Work
One of the major review datasets was released by Amazon, Inc. back in 2020 Keung et al. (2020). As an initiative separate from the Amazon, Hou et al. (2024) released up-to-date Amazon review datasets named “AMAZON REVIEWS 2023”, ranging from May 1996 to September 2023  [^8] https://amazon-reviews-2023.github.io. There are user-generated review datasets in the hotel domain, mainly crawled from leading online travel platforms such as TripAdvisor and Booking.com. For instance, Fang et al. (2016) crawled 41k reviews posted in 2014 for attractions in New Orleans to analyze the perceived helpfulness of reviews. Tsamis et al. (2021) crawled 65k English hotel reviews from TripAdvisor and trained a DNN on helpfulness prediction. The Kaggle dataset  [^9] https://www.kaggle.com/datasets/jiashenliu/515k-hotel-reviews-data-in-europe/discussion?sort=undefined contains 515k reviews crawled from Booking.com published between August 2015 and August 2017. Finally, Igebaria et al. (2024) created a corpus of English reviews posted in 2023 for predicting individual review helpfulness. Although these datasets are large-scale, they do not provide sufficiently long period, covering three significant eras in the Travel industry: pre-COVID, mid-COVID, and post-COVID. None of these studies focuses on releasing review texts for more than 15 years in the hotel domain.
@@ 4 Factors Causing Data Drift
“エモい” is originated from “emotional”. This is a very convenient and ambiguous term that encompasses a diverse range of emotions and sensations such as "feeling deeply moved," “resonating deeply,“nostalgic,“poignant, and “having a unique atmosphere. Additionally, although there was no significant difference, other new slang words appeared in 2024 whereas in 2019 these did not such as “チルい derived from “Chilling out” and “沼る (obsessed)”. Such new slang words make NLP tasks, such as sentiment analysis, more difficult.